# Accurate Screening of Functional Materials with Machine-Learning Potential and Transfer-Learned Regressions: Heusler Alloy Benchmark

### v1

@@ 2.6 Frozen transfer learning for MLRM construction
To improve the performance of the MLRM, we employed a frozen transfer learning strategy using the eSEN-30M-OAM MLIP as the base model. The eSEN-30M-OAM MLIP was trained on the OMat, MPtrj, and sAlexandria databases, providing a comprehensive training set spanning the periodic table [13, 32, 23, 33, 34]. Through this training, the embedding and the first several layers learn general chemical and structure representations. To leverage this, we transferred the parameters of the embedding and the first several layers into our MLRM and kept them fixed (frozen layers), while only updating the remaining layers and the output layer (flexible layers). This approach is analogous to a recent work in which the initial layers of ORB, EqV2, or MACE MLIPs were used to generate feature vectors that were subsequently passed to property prediction models such as MODNet, XGBoost, and MLP [35].
@@
The eSEN-30M-OAM model consists of 10 layers. Figure 5 shows the $R^{2}$ scores of models trained with different numbers of frozen layers, denoted TL-MLIP-$n$. In the $n=0$ case, the embedding layer is also flexible. Results for $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$ are presented. Model performance improves as the number of frozen layers increases up to $n=7$, after which it declines when more layers are frozen. This trend reflects the balance between transferring knowledge from the base model and maintaining sufficient flexibility to adapt to the new task. Based on this analysis, TL-MLIP-7 was selected as the final model for $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, $E_{\mathrm{aniso}}$, and $\{\boldsymbol{m}_{i}\}$ MLRM. For comparison, a model trained from scratch (w/o TL) is included, which yields lower $R^{2}$ scores and highlights the benefit of transfer learning. We also tested a variant initialized from a base model pre-trained on formation energy ($\Delta E$) data from the HeuslerDB database (denoted TL-$\Delta E$-0). TL-MLIP-0 outperforms TL-$\Delta E$-0, indicating that initialization from the eSEN-30m-oam model, trained on a comprehensive dataset, is more advantageous.
@@ 3 Discussion
We demonstrated the feasibility of combining MLIP s and MLRM s for HTP screening. As a case study, we identified 334 conventional quaternary Heusler compounds and 924 all-$d$ Heusler compounds that exhibit thermodynamic, dynamical, and magnetic stability, together with large $E_{\mathrm{aniso}}$. The precision of this workflow was confirmed through DFT validation of the candidate list.
When a database for a subset of materials is available, the same workflow can be applied to systematically explore the remaining chemical space. For novel materials, MLIP can first be used to reduce the search space by filtering for thermodynamic stability, while target properties are computed for a representative subset of compounds to train MLRMs for ML-HTP screening of the broader space. The workflow can also be applied iteratively, with results from DFT validations fed back into the ML models to improve their accuracy and predictive power, thereby generating new candidate materials and recovering promising compounds that may have been missed in earlier screening stages.
In the framework of this study, MLIP perform the critical task of structure optimization, which is traditionally handled by DFT-derived methods, while MLRMs predict the properties of the optimized structures, tasks that are typically carried out using DFT or DFT-derived methods. Together, MLIP and MLRMs enable a drop-in replacement for DFT in conventional HTP workflows. This replacement is general and can be readily extended to other properties, material classes, and DFT-based HTP pipelines, enabling accelerated HTP screening and discovery of new materials.
@@ 4 Methods
4.1 Composition and structure
This study screened two classes of Heusler compounds not included in HeuslerDB: (i) conventional quaternary Heuslers and (ii) all-$d$ ternary and quaternary Heuslers. The ternary Heuslers were considered in regular, inverse, and half-Heusler structures [20]. For conventional quaternary compounds, we enumerated all combinations where $X$ and $Y$ are transition metals from the $d$-block (excluding Tc and Hg), and $Z$ is a main-group element from groups 13, 14, or 15 of the $p$-block, as illustrated in Fig. 1(b). In addition, La and Lu were included for $X$ and $Y$ because their 4ff orbitals are either empty or fully filled. This exhaustive enumeration, accounting for symmetry constraints, yielded 131,544 unique compositions. For the all-$d$ Heuslers, we extended the combinatorial space to include only $d$-block transition metals together with La and Lu across all four sites ($X_{1}$, $X_{2}$, $Y$, and $Z$), resulting in a separate set of 104,139 unique compositions.
@@ 4.2 ML-HTP workflow
The ML-HTP workflow, including the use of MLIP and MLRM together with the selection criteria, is schematically illustrated in Fig. 1(c) and described in detail below.
(a) Structure optimization
The initial lattice constant was estimated as the average value of known $X_{2}YZ$ Heusler compounds in the HeuslerDB database that share two elemental species with the target composition. A conventional cell in cubic phase with this estimated lattice constant was then constructed. To generate initial structures, the lattice parameters of this cell were systematically varied: the $a$, $b$, and $c$ were uniformly scaled by $\pm 10\%$ or $\pm 30\%$, and, alternatively, the $c$ alone was scaled by $\pm 10\%$ or $\pm 30\%$. All generated structures were subsequently converted to the primitive unit cell and relaxed using the eSEN-30M-OAM MLIP. The structure with the lowest energy after relaxation was identified as the ground state. The choice of eSEN-30M-OAM was motivated by its superior performance relative to other MLIPs, as discussed in Sec. 2.3. The selection of initial structures was validated by a convergence test, which is provided in the Supplementary Materials.
(b) Thermodynamic stability from MLIP energies
For a compound to be thermodynamically stable against decomposition into its constituent elements or competing phases, the formation energy must be negative ($\Delta E<0$) and the distance to the convex hull must be zero ($\Delta H=0$). In practice, however, metastable phases ($\Delta H>0$) at 0 K may become stabilized at finite temperature [36]. Following our previous work, we adopt a practical stability criterion of $\Delta E<0.0$ eV/atom and $\Delta H<0.22$ eV/atom, which has been shown to effectively capture experimentally accessible compounds [20]. Using the energies of ground state candidates, constituent elements and competing phases predicted by eSEN-30M-OAM, we computed $\Delta E$ and $\Delta H$ to assess thermodynamic stability.
(c) Property prediction using MLRMs
@@
(d) DFT validation
The candidate list was validated with DFT calculations to assess the reliability of the ML-HTP workflow. Structure optimizations were performed starting from various initial spin configurations, consistent with our previous DFT-THP work. For conventional Heusler compounds, the magnetic moments at the $X_{1}$, $X_{2}$, and $Y$ sites were initialized in configurations where they were either parallel or antiparallel to each other. For all-$d$ Heuslers, spin arrangements on all four sites were considered. To capture possible high-spin and low-spin states, two initial magnitudes of the local moments ($|\boldsymbol{m}_{i}|=1$ and $4~\mu_{\mathrm{B}}$) were tested, along with a non-magnetic configuration ($|\boldsymbol{m}_{i}|=0$). The MLIP-optimized structures served as the starting geometries. After structure relaxation, the ground state was identified by comparing total energies. For the resulting ground states, we computed $\Delta E$, $\Delta H$, $E_{\mathrm{aniso}}$, phonon, and $T_{{\rm c}}$ using VASP, OQMD, ALAMODE, and SPRKKR [37, 38, 39, 40, 41, 42].
@@ 4.3 Computational methods
The MLIP-based structural optimizations were performed using the Atomic Simulation Environment (ASE) package [43]. The Fast Inertial Relaxation Engine (FIRE) optimizer was employed, with symmetry constraints enforced throughout the relaxation process [44]. To ensure consistency in reference energies for computing $\Delta E$ and $\Delta H$, the elemental phases and competing phases were also optimized using the same MLIP. Their initial geometries were taken from DFT-optimized structures in the Open Quantum Materials Database (OQMD) database [37, 38].
@@
The MLRMs were developed to predict $\omega_{\mathrm{min}}$, $\{\boldsymbol{m}_{i}\}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$. To leverage prior knowledge, we employed a frozen transfer learning strategy, as illustrated in Fig. 1(a). Each MLRM was initialized from the pretrained eSEN-30M-OAM MLIP, with the embedding layer and the first seven message-passing layers kept frozen, while the final three layers and the output layer were fine-tuned. This framework was implemented using a modified version of the FAIRChem package (v1) [45]. For the $\{\boldsymbol{m}_{i}\}$ MLRM training, the loss function was adapted to address the global sign ambiguity of magnetic moments, as described in Sec. 2.5. The training dataset consisted of HeuslerDB together with newly computed $E_{\mathrm{aniso}}$ and $T_{{\rm c}}$ values based on the optimized structures in HeuslerDB. The data were randomly partitioned into training, validation, and test sets in an 8:1:1 ratio. For $\{\boldsymbol{m}_{i}\}$, we used all ground-state entries, resulting in 27,864 data points. The $\omega_{\mathrm{min}}$ data were available for all thermodynamically stable ground states, yielding 8,198 entries. For $T_{{\rm c}}$, 2,106 data points were used, including 750 newly computed values. Since $E_{\mathrm{aniso}}$ data were not included in HeuslerDB, we calculated $E_{\mathrm{aniso}}$ for all magnetic tetragonal ground-state systems, obtaining 6,123 entries.
@@
DFT calculations were performed primarily with Vienna ab initio Simulation Package (VASP) [46, 47], using the projector augmented wave (PAW) method and the generalized gradient approximation (GGA) with the Perdew–Burke–Ernzerhof (PBE) functional [48, 49]. $\Delta E$, $\Delta H$, phonon, and $T_{{\rm c}}$ were computed using OQMD, ALAMODE, and SPRKKR following the methodology of our previous DFT-HTP study of ternary Heusler compounds [37, 38, 39, 40, 41, 42, 20]. Phonon calculations for the all-$d$ compounds MnOsMnRe and MnReMnRu failed due to convergence issues in DFT, and these compounds were treated as unstable in the validation rate analysis. $E_{\mathrm{aniso}}$ was calculated as $_{\mathrm{aniso}} =EE$$_{\perp}-E_{\|}$ using the force theorem [50, 51, 52]. Calculations were performed in the primitive cell with $\boldsymbol{k}$-meshes generated using Python Materials Genomics (pymatgen) at a density of $6000\text{\,}\mathrm{\text{Å}}^{-3}$, and the tetrahedron method with Blöchl corrections was applied [53, 54]. Input generation, structure manipulation, and symmetry analysis were carried out using pymatgen, ASE, ASE2SPRKKR, and spglib [53, 43, 55, 38, 56].

### v2

@@ 2.6 Frozen transfer learning for MLRM construction
To improve the performance of the MLRM, we employed a frozen transfer learning strategy using the eSEN-30M-OAM uMLIP as the base model. The eSEN-30M-OAM uMLIP was trained on the OMat, MPtrj, and sAlexandria databases, providing a comprehensive training set spanning the periodic table [barroso-luqueOpenMaterials20242024, jainCommentaryMaterialsProject2013, dengCHGNetPretrainedUniversal2023, schmidtMachineLearningAssistedDeterminationGlobal2023, schmidtImprovingMachinelearningModels2024]. Through this training, the embedding and the first several layers learn general chemical and structure representations. To leverage this, we transferred the parameters of the embedding and the first several layers into our MLRM and kept them fixed (frozen layers), while only updating the remaining layers and the output layer (flexible layers). This approach is analogous to a recent work in which the initial layers of ORB, EqV2, or MACE uMLIPs were used to generate feature vectors that were subsequently passed to property prediction models such as MODNet, XGBoost, and MLP [kimLeveragingNeuralNetwork2025a].
@@
The eSEN-30M-OAM model consists of 10 layers. Fig. 6 (a) shows the $R^{2}$ scores of models trained with different numbers of frozen layers, denoted TL-uMLIP-$n$. In the $n=0$ case, the embedding layer is also flexible. Results for $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$ are presented. Model performance improves as the number of frozen layers increases up to $n=7$, after which it declines when more layers are frozen. This trend reflects the balance between transferring knowledge from the base model and maintaining sufficient flexibility to adapt to the new task. Based on this analysis, TL-uMLIP-7 was used in the ML-HTP case study for $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, $E_{\mathrm{aniso}}$, and $\{\boldsymbol{m}_{i}\}$ MLRM. For comparison, a model trained from scratch (w/o TL) is included, which yields lower $R^{2}$ scores and highlights the benefit of transfer learning. We also tested a variant initialized from a base model pre-trained on formation energy ($\Delta E$) data from the HeuslerDB database (denoted TL-$\Delta E$-0). TL-uMLIP-0 outperforms TL-$\Delta E$-0, indicating that initialization from the eSEN-30M-OAM model, trained on a comprehensive dataset, is more advantageous.
Figure 6: Frozen transfer learning performance and generalization. a) $R^{2}$ of models initialized from the eSEN-30M-OAM uMLIP and trained with different numbers of frozen layers (denoted TL-uMLIP-$n$). In the $n=0$ case, the embedding layer is also trainable. A model trained from scratch (denoted w/o TL) and a transfer learning variant initialized from a base model pre-trained on $\Delta E$ data from the HeuslerDB database (denoted TL-$\Delta E$-0) are included for comparison. b) Comparison of model performance on group-wise splits versus random splits. Models were evaluated on predicting $T_{{\rm c}}$. Results are shown for frozen transfer learning (FTL) and models trained from scratch (w/o TL). For each test, three neighboring elements within the same period were used as the holdout elements. Six sets of d-block elements and four sets of p-block elements were tested.
Besides improving overall performance, transfer learning from uMLIP can significantly enhance model generalization to unseen elements. To demonstrate this, we evaluated model performance using group-wise split analysis. For each group-wise split test, three neighboring elements within the same period were used as the holdout elements. Any compound containing at least one of the holdout elements was reserved as the holdout test data; the remaining compounds were split into train/val with a 9:1 ratio. This setup simulates realistic scenarios where models are applied to material systems containing elements not present in the training data, a typical domain shift in materials science. For direct comparison, we also measured performance using train/val/test sets created using random splits while keeping the counts of train/val/test the same as in the group-wise split. We tested six sets of d-block elements and four sets of p-block elements as holdout sets. We used $T_{{\rm c}}$ as the target property for efficiency. Results obtained using frozen transfer learning and models trained from scratch are shown in Fig. 6 (b).
Using FTL, performance on group-wise splits is generally lower than on random splits, indicating that the model performance on unseen elements is reduced. The drop in performance for p-block holdouts is smaller, which is consistent with the fact that p-block elements are typically nonmagnetic in these compounds and thus less directly related to the target property $T_{{\rm c}}$. For group-wise split tests with models trained from scratch, a much larger performance decrease is observed. This occurs because FTL transfers the general chemical representation learned by uMLIP to property regressors, whereas models trained from scratch do not benefit from this prior knowledge. This highlights the advantage of transfer learning from uMLIP for improving generalization to unseen elements. Variation in random-split performance is due to different train set sizes since the number of holdout compounds varies with the chosen holdout elements.
It should be noted that TL is not always effective. It can be limited when the representations learned by base model do not generalize to the target domain. For example, source and target may differ substantially in material class (e.g. bulk versus surface or molecular systems), or downstream task concerns different physics (e.g. static energetics versus dynamic properties) [yamadaPredictingMaterialsProperties2019a, chenGraphNetworksUniversal2019a, changOvercomingDataScarcity2022]. Furthermore, a recent work, which transfer from model trained using computational data to learn experimental data, revealed that the error of the transfer-learned model decreases according to a power-law as the size of the computational data for the base model increases [minamiScalingLawSim2Real2025]. This suggests that transfer learning may also be less effective when the base model is trained on a small dataset even if the base model performance within its own domain is good.
@@ 3 Discussion
We demonstrated the feasibility of combining uMLIP s and MLRM s for HTP screening. As a case study, we identified 334 conventional quaternary Heusler compounds and 924 all-$d$ Heusler compounds that exhibit thermodynamic, dynamical, and magnetic stability, together with large $E_{\mathrm{aniso}}$. The precision of this workflow was confirmed through DFT validation of the candidate list.
For other material systems, if database for a subset is available, the same workflow can be applied to explore the remaining chemical space. For novel materials, uMLIP can be used directly to reduce the search space by filtering for thermodynamic stability, while target properties need to be computed for a representative subset of compounds to train MLRMs for ML-HTP screening of the broader space.
The MLRM s in this study involve a domain shift, as the train dataset does not contain quaternary or all-$d$ Heusler compounds, yet the models are applied to these systems. While performance is satisfactory, further gains are possible using active learning approach. Iteratively refining the models by selecting informative compounds can improve performance and help identify candidates missed by the current one-shot approach. In such iterative framework, current DFT validation results can be added to the training set, forming the first iteration of model refinement. This iterative strategy naturally extends the present work, enabling more thorough exploration of chemical space while progressively improving model accuracy. Such an approach is especially valuable when the target material space differs substantially from the training data.
In the framework of this study, uMLIP perform the critical task of structure optimization, which is traditionally handled by DFT-derived methods, while MLRMs predict the properties of the optimized structures, tasks that are typically carried out using DFT or DFT-derived methods. Together, uMLIP and MLRMs enable a drop-in replacement for DFT in conventional HTP workflows. This replacement is general and can be readily extended to other properties, material classes, and DFT-based HTP pipelines, enabling accelerated HTP screening and discovery of new materials.
@@ 4.1 ML-HTP workflow
The ML-HTP workflow is schematically illustrated in Fig. 1(b) and described in detail below.
In step (a) of ML-HTP workflow, the structures were optimized and the $\Delta E$ and $\Delta H$ were calculated using uMLIP. The initial lattice constant was estimated as the average value of known $X_{2}YZ$ Heusler compounds in the HeuslerDB database that share two elemental species with the target composition. A conventional cell in cubic phase with this estimated lattice constant was then constructed. To generate initial structures, the lattice parameters of this cell were systematically varied: the $a$, $b$, and $c$ were uniformly scaled by $\pm 10\%$ or $\pm 30\%$, and, alternatively, the $c$ alone was scaled by $\pm 10\%$ or $\pm 30\%$. All generated structures were subsequently converted to the primitive unit cell and relaxed using the eSEN-30M-OAM uMLIP. The structure with the lowest energy after relaxation was identified as the ground state. The choice of eSEN-30M-OAM was motivated by its superior performance relative to other uMLIPs, as discussed in Sec. 2.3. The selection of initial structures was validated by a convergence test, which is provided in Fig. S4.
For a compound to be thermodynamically stable against decomposition into its constituent elements or competing phases, the formation energy must be negative ($\Delta E<0$) and the distance to the convex hull must be zero ($\Delta H=0$). In practice, however, metastable phases ($\Delta H>0$) at 0 K may become stabilized at finite temperature [sunThermodynamicScaleInorganic2016]. Following our previous work, we adopt a practical stability criterion of $\Delta E<0.0$ eV/atom and $\Delta H<0.22$ eV/atom, which has been shown to effectively capture experimentally accessible compounds [xiaoHighthroughputComputationalScreening2025]. Using the energies of ground state candidates, constituent elements and competing phases predicted by eSEN-30M-OAM, we computed $\Delta E$ and $\Delta H$ to assess thermodynamic stability.
@@
In step (c) of workflow, the candidate list was validated with DFT calculations to assess the reliability of the ML-HTP workflow. Structure optimizations were performed starting from various initial spin configurations, consistent with our previous DFT-HTP work. For conventional Heusler compounds, the magnetic moments at the $X_{1}$, $X_{2}$, and $Y$ sites were initialized in configurations where they were either parallel or antiparallel to each other. For all-$d$ Heuslers, spin arrangements on all four sites were considered. To capture possible high-spin and low-spin states, two initial magnitudes of the local moments ($|\boldsymbol{m}_{i}|=1$ and $4~\mu_{\mathrm{B}}$) were tested, along with a nonmagnetic configuration ($|\boldsymbol{m}_{i}|=0$). The uMLIP-optimized structures served as the starting geometries. After structure relaxation, the ground state was identified by comparing total energies. For the resulting ground states, we computed $\Delta E$, $\Delta H$, $E_{\mathrm{aniso}}$, phonon, and $T_{{\rm c}}$ using VASP, OQMD, ALAMODE, and SPRKKR [saalMaterialsDesignDiscovery2013, bahnObjectorientedScriptingInterface2002, tadanoSelfconsistentPhononCalculations2015, AnharmonicForceConstants, ebertCalculatingCondensedMatter2011, liechtensteinLocalSpinDensity1987].
@@ 4.2 Computational methods
The uMLIP-based structural optimizations were performed using the Atomic Simulation Environment (ASE) package [hjorthlarsenAtomicSimulationEnvironment2017]. The Fast Inertial Relaxation Engine (FIRE) optimizer was employed, with symmetry constraints enforced throughout the relaxation process [bitzekStructuralRelaxationMade2006]. To ensure consistency in reference energies for computing $\Delta E$ and $\Delta H$, the elemental phases and competing phases were also optimized using the same uMLIP. Their initial geometries were taken from DFT-optimized structures in the Open Quantum Materials Database (OQMD) database [saalMaterialsDesignDiscovery2013, bahnObjectorientedScriptingInterface2002].
@@
The MLRMs were developed to predict $\omega_{\mathrm{min}}$, $\{\boldsymbol{m}_{i}\}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$. To leverage prior knowledge, we employed a frozen transfer learning strategy, as illustrated in Fig. 1(a). Each MLRM was initialized from the pretrained eSEN-30M-OAM uMLIP, with the embedding layer and the first seven message-passing layers kept frozen, while the final three layers and the output layer were fine-tuned. This framework was implemented using a modified version of the FAIRChem package (v1) [shuaibiFAIRChem2025]. For the $\{\boldsymbol{m}_{i}\}$ MLRM training, the loss function was adapted to address the global sign ambiguity of magnetic moments, as described in Sec. 2.5. The training dataset consisted of HeuslerDB together with newly computed $E_{\mathrm{aniso}}$ and $T_{{\rm c}}$ values based on the optimized structures in HeuslerDB. The data were randomly partitioned into training, validation, and test sets in an 8:1:1 ratio. For $\{\boldsymbol{m}_{i}\}$, we used all ground-state entries, resulting in 27,864 data points. The $\omega_{\mathrm{min}}$ data were available for all thermodynamically stable ground states, yielding 8,198 entries. For $T_{{\rm c}}$, 2,106 data points were used, including 750 newly computed values. Since $E_{\mathrm{aniso}}$ data were not included in HeuslerDB, we calculated $E_{\mathrm{aniso}}$ for all magnetic tetragonal ground-state systems, obtaining 6,123 entries.
@@
DFT calculations were performed primarily with Vienna ab initio Simulation Package (VASP) [kresseEfficiencyInitioTotal1996, kresseEfficientIterativeSchemes1996], using the projector augmented wave (PAW) method and the generalized gradient approximation (GGA) with the Perdew–Burke–Ernzerhof (PBE) functional [kresseUltrasoftPseudopotentialsProjector1999, perdewGeneralizedGradientApproximation1996]. $\Delta E$, $\Delta H$, phonon, and $T_{{\rm c}}$ were computed using OQMD, ALAMODE, and SPRKKR following the methodology of our previous DFT-HTP study of ternary Heusler compounds [saalMaterialsDesignDiscovery2013, bahnObjectorientedScriptingInterface2002, tadanoSelfconsistentPhononCalculations2015, AnharmonicForceConstants, ebertCalculatingCondensedMatter2011, liechtensteinLocalSpinDensity1987, xiaoHighthroughputComputationalScreening2025]. The agreement of magnetic moments between VASP and SPRKKR is demonstrated in Fig. S7. Phonon calculations for the all-$d$ compounds MnOsMnRe and MnReMnRu failed due to convergence issues in DFT, and these compounds were treated as unstable in the validation rate analysis. $E_{\mathrm{aniso}}$ was calculated as $_{\mathrm{aniso}} =EE$$_{\perp}-E_{\|}$ using the force theorem [daalderopFirstprinciplesCalculationMagnetocrystalline1990, xingLatticeDynamicsIts2022, xingFirstprinciplesPredictionPhase2023]. Calculations were performed in the primitive cell with $\boldsymbol{k}$-meshes generated using Python Materials Genomics (pymatgen) at a density of $6000\text{\,}\mathrm{\text{Å}}^{-3}$, and the tetrahedron method with Blöchl corrections was applied [ongPythonMaterialsGenomics2013, blochlImprovedTetrahedronMethod1994]. Input generation, structure manipulation, and symmetry analysis were carried out using pymatgen, ASE, ASE2SPRKKR, and spglib [ongPythonMaterialsGenomics2013, hjorthlarsenAtomicSimulationEnvironment2017, ASE2SPRKKRSoftwarePackage, bahnObjectorientedScriptingInterface2002, togoSpglibSoftwareLibrary2024].
4.3 Computational cost
The computational cost for each task in the HTP workflow is summarized in Table 2. DFT timings are based on validation runs for ML-selected candidates, while ML timings are taken from the ML-HTP screening. For DFT, we report per-job statistics such as mean, median, and interquartile range (IQR). uMLIP structure optimization and MLRM predictions were performed in batches, and mean values are reported as individual per-job timings are not available. MLRM training times correspond to the wall time required to train a single model on one GPU. DFT calculations except phonon were performed on dual-socket Intel Xeon Platinum 8268 (Cascade Lake, 24 cores per CPU, 2.9 GHz, 48 cores per node), phonon calculations were performed on Fujitsu A64FX processors (Armv8.2-A SVE 512 bit, 48 compute cores per CPU, 2.2 GHz), uMLIP structure optimizations were performed on dual-socket Intel Xeon Gold 6230 (Cascade Lake, 20 cores per CPU, 2.1 GHz, 40 cores per node), and MLRMs were trained and applied on NVIDIA RTX 6000 Ada GPUs. All DFT calculations, except for phonon calculations, were performed on 2 nodes; phonon calculations were performed on 6 nodes. The uMLIP structure optimizations were performed on 1 node and MLRM training and prediction were performed on 1 GPU.
DFT (mean/median/IQR) ML training (mean) ML prediction (mean)
Structure relaxation 5.6/3.8/3.0 0.12
$\{\boldsymbol{m}_{i}\}$ 876 <<1e-03
Phonon stability 4542/4008/1442 258 <<1e-03
$T_{{\rm c}}$ 70/66/14 78 <<1e-03
$E_{\mathrm{aniso}}$ 142/133/34 192 <<1e-03
Table 2: Computational cost of each task in the HTP workflow, reported in node-minutes. DFT entries show per-job statistics (mean/median/IQR). ML entries show mean values. For structure relaxation, one job corresponds to a single relaxation from an initial distortion. For structure optimization, the cost per compound should be multiplied by the number of initial structures, which varies across material systems; here, we report the cost for a single structure relaxation. The uMLIP is pre-trained and used without fine-tuning, so its training time is excluded. DFT structure optimizations directly yield local magnetic moments, so no separate timing is reported.
In the ML‑HTP case study, uMLIP structure optimization, MLRM training, and MLRM prediction used 1,835 node‑hours, 23 GPU‑hours, and 4 GPU‑hours, respectively. DFT calculations of ML‑selected candidates consumed 1,645 node‑hours for optimizations and 99,680 node‑hours for property evaluations. Ignoring the node difference, the total ML‑HTP cost is 103,160 node‑hours plus 27 GPU‑hours. A DFT-HTP workflow screening the same chemical space would need about 256,160 node‑hours for structure optimizations and 18,802,625 node‑hours for property evaluations, estimated using statistics reported in Table 2. This comparison shows an estimated speed‑up of 185 times for ML‑HTP with DFT validation, or 104 times if used without DFT validation at the end. Please note that these estimates are approximate, and waiting time for job execution, human time spent on debugging and workflow management are not included. The actual speed-ups also vary based on material systems, target properties, and computational resources.